OpenStudy (dls):

If g(x) = ..... and x=0 is a part of maxima of y=f(x) then

OpenStudy (dls):

1)x=0 is max for g(x) 2)x=0 is min for g(x) 3)x=0 is inflexion for g(x) 4)Nothing can be said.

OpenStudy (dls):

\[\Huge g(x)=3f(x)^2-2f(x)^3-6f(x)\]

OpenStudy (dls):

@amistre64 @dan815 @terenzreignz

terenzreignz (terenzreignz):

something tells me it's the 4th... way too many possibilities...

OpenStudy (dls):

umm no..

terenzreignz (terenzreignz):

It isn't? interesting :D

OpenStudy (dls):

@hartnn

terenzreignz (terenzreignz):

Let's just differentiate, I suppose :3 \[\large g'(x) =\left[f(x) -f(x)^2-1\right]f'(x)\]

OpenStudy (dls):

yup i got that,we also have 6 outside

terenzreignz (terenzreignz):

Whoops \[\large g'(x) =6\left[f(x) -f(x)^2-1\right]f'(x)\]

terenzreignz (terenzreignz):

And differentiate again? (this'll be a nightmare)

terenzreignz (terenzreignz):

\[\Large \frac{g''(x)}6 =\left[1-2f(x)\right]\left[f'(x)\right]^2 + \left[f(x) -f(x)^2-1\right]f''(x) \]

OpenStudy (dls):

..

OpenStudy (dls):

well if you observe if f'(x) that thing's D is <0..

terenzreignz (terenzreignz):

Well, obviously, g'(0) is 0, so it's a possible extrema

terenzreignz (terenzreignz):

because f'(0) is 0

terenzreignz (terenzreignz):

So we need g''(0)

terenzreignz (terenzreignz):

\[\Large \frac{g''(0)}6 =\left[1-2f(0)\right]\left[f'(0)\right]^2 + \left[f(0) -f(0)^2-1\right]f''(0)\]

terenzreignz (terenzreignz):

This part is zero (since f'(0) is zero) \[\Large \frac{g''(0)}6 =\color{red}{\left[1-2f(0)\right]\left[f'(0)\right]^2 }+ \left[f(0) -f(0)^2-1\right]f''(0)\]

terenzreignz (terenzreignz):

So... \[\Large \frac{g''(0)}6 =\left[f(0) -f(0)^2-1\right]f''(0)\]

terenzreignz (terenzreignz):

This is negative \[\Large \frac{g''(0)}6 =\left[f(0) -f(0)^2-1\right]\color{blue}{f''(0)}\] since 0 is a maximum of f, second derivative at that point is negative...

terenzreignz (terenzreignz):

And this part \[\Large \frac{g''(0)}6 =\color{blue}{\left[f(0) -f(0)^2-1\right]}f''(0)\] is also negative, for whatever f(0) is... You can check the graph of \[\Large x - x^2 - 1\] It's all negative for whatever x is.

terenzreignz (terenzreignz):

Now, that means \[\Large \frac{g''(0)}{6}\] is positive meaning \[\Large g''(0)\] is positive....

terenzreignz (terenzreignz):

\[\Large g'(0)=0\]\[\Large g''(0)>0\] Means 0 is a minimum of g Case closed :P